An investor owns 1,000 shares of BXO shares and 500 shares of MBA stock. BXO shares are currently selling for $23 each and have a beta of 1.4. MBA shares are currently selling for $55 each and have a beta of 1.2. What is the beta of the portfolio of shares owned by this investor?

____

A)        1.20

B)        1.23

C)        1.29

D)        1.33
Answer #1


Solution: BXO MBA Shares Price investment Amount Weight Beta weighted beta b c =a*b d=c/50500 f=de 1000 23 23000 45.54% 1.4 0